OpenStudy (deadshot):

The length of the hypotenuse of a 30°-60°-90° triangle is 14 units. Find the exact value of the length of the short leg.

OpenStudy (mathstudent55):



OpenStudy (deadshot):

so would it be 14/cos60?

OpenStudy (deadshot):

or 14/sec60?

OpenStudy (mathstudent55):

14 * cos 60

OpenStudy (deadshot):

oh

OpenStudy (mathstudent55):

cos 60 = a/14 a = 14 * cos 60

OpenStudy (deadshot):

i get 0.5 for cos60

OpenStudy (deadshot):

so then 14*cos60=7?

OpenStudy (deadshot):

so the answer would be 7 units, right?

OpenStudy (deadshot):

Thanks!